1 2 -1 -1 | if it exists. Check your answer. Find the inverse of the matrix | 1 1 3 -1 1

show all steps thanks
Answer #1

12 IT Let A= ( 1- 2. 4 podo uln Buipwindseg to first row) Then det (A)= 1 이지혜세혜 4(1-1)-2(+5)-1(-3) = -8 +4 = = 4 0 As determi